From nch@roe.ac.uk Wed Sep 29 13:11:29 2010 Date: Fri, 24 Sep 2010 15:53:34 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren , Tom Kerr Subject: WFAU VDFS Science Archive project meet mins, 24/09/10 Minutes of WFAU VDFS Science Archive meeting: September 24th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NJC, NCH, RSC, ETWS Apologies: JPE, AL, RGM, KTN, MSH, MAR, RPB DoNM: Friday October 8th in the Vista Hut Actions discharged this week: ----------------------------- ACTION: NCH to assign actions on the following next time: - Modify entire code base to work with Tiles (Partly discharged - but needs fully testing on the new V1 data). This includes checking code works here on deep stacks, checking quality control/ QBF, ingest, CU7, CU6. Also do we need to do better links between tile detections and pawprint detections or will Provenance etc be good enough. (RSC has added tile qbits; also suggested a preliminary test run of CU7 using shallow tiles first of all). - VVV SynopticSource (allowing both SourceXSynopticSourceBestMatch and SourceXDetectionBestMarch) Discharged thanks to much work by all, esp. NJC Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to remember to mention SSA crossneighbour production in WSA/VSA next week (it seems to be switched off?) Discharged; *SourceXSSASource are absent from UKIDSSDR4 onwards while they are present in VSA release DBs. Someone to sort out (maybe by capitalising the create table statements in relevant scripts?) Actions carried forward from 10/09/10 meeting --------------------------------------------- ACTION: TBD to draft and MOU on behalf of PSPIs for ESO-SAF deliveries Continues; on hold until at least next PSPI teleconf, and possibly until "Phase 3" workshop at ESO. ACTION: NJC to suggest workshop at next PSPI meeting. NJC noted that he needs to liaise with JPE about the next teleconf. Specific points and new actions: -------------------------------- Project management: Nothing new this week. WFCAM & VISTA updates: No updates this week. Comments and issues arising from CASU minutes: No new minutes this week. Networking: WFCAM 10B data started transfering this week following restoration of site power (see below). CASU have notified us of a partial reprocessing of the VIRCam SV data (v1.1?!) nights, so a small amount of early VVV/VMC is presumably included in this (all agreed that this is lower priority at present). WSA/VSA Operations: ETWS noted -- VISTA catalogues have been ingested up to 23/07/2010 for all surveys but VVV. -- The VVV data will be ingested once the seperate VVV database is created to enable easier ingest and release curation for the other surveys. -- The latest WFCAM 10B data (August 2010) has been transferred and metadata ingests will be completed at the and of today. -- Enhanced the DataDeamon output to be more informative and more distintive between VSA and WSA curation. Hardware and Systems: Major news this time is that on Thursday 16th September at 5am there was a site-wide power failure that was not corrected until Tuesday 21st. This caused huge disruption to curation activities, and also inconvenienced many users. Web documentation has been updated by NJC to note the loss of service; much email support requests where handled by NCH and MAR sent out apologies and all-clear to the announce lists. NJC noted that the need for a new beefy 64bit curation server is getting urgent, as scale-out testing of deep stacking for the VSA is clashing with operations. ETWS noted that khafre may have some memory issues since it's update to the latest Debian release, with errors being reported from the "north-bridge". RPB will go looking for trolls ... Software: RSC noted that there will be undoubtedly lots of small mods required to various bits of the curation SW to cope with splitting off the humungous VVV into it's own ingest DB. Otherwise, nothing else of note this week. Survey Data Release: UDS and VIDEO mosaic deliveries are pending for UKIDSS DR8 and VIDEO v1.0 releases. NJC is reviewing the TWiki documentation for the former so that operations can incorporate the UDS into DR8. Non-survey Data Release: Nothing of note this week. Astrogrid deployment & Data Analysis services: No news this week. Miscellaneous: ETWS and KTN are off to Austria... sorry, southern Germany next week to attend Oktoberfest... sorry, the VST "readiness review" meeting in Garching near Munich. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================